OpenStudy (anonymous):

solve the equation 2^3x=16 please help! thx

OpenStudy (anonymous):

2^4 = 16 => 3x=16 => x= 16/3

OpenStudy (campbell_st):

nearly... 2^4 = 16 so 3x = 4 equating powers... then x = 4/3

OpenStudy (anonymous):

where did you get ^4

OpenStudy (campbell_st):

16 = 2^4 so 2 x 2 x 2 x 2 = 16

OpenStudy (anonymous):

thx!:-)